# Tracking Any Object Amodally

### v2

@@ 3.1 Statistical Analysis
We compare the statistics of TAO-Amodal to other amodal benchmarks in Table 1. For NuScenes, which only categorizes object visibilities into four buckets, we use interpolation to estimate the number of boxes below visibility 0.1 and 0.8. A few amodal datasets are omitted from the table either because they have been incorporated in TAO-Amodal [7, 61] or because these datasets lack quantified visibilities for categorizing distinct occlusion scenarios [9, 51]. TAO-Amodal encompasses annotations across an extensive 880 categories, which can be used to learn and evaluate object priors in a large-vocabulary setting, in contrast to benchmarks that only cover people or vehicles. Additionally, TAO-Amodal features a 10$\times$ longer evaluation duration, ensuring its status as a challenging benchmark.
Figure 3: Amodal Expander Architecture. Given $N$ flattened proposal features and modal box (delta) predictions represented with 256-dim positional encodings [56], we predict amodal box (deltas) with a two-layer MLP (unless otherwise specified). Further architecture details could be found in Sec. 5.2
@@
Table 2: Off-the-shelf trackers on TAO-Amodal validation set. We use the metrics defined in Sec. 5.1, where the visibility range is indicated by the superscript. Off-the-shelf trackers were either trained on TAO [10] or on synthetic videos [64] generated using LVIS images [21], with categories aligned with our dataset. We use visibility attributes to analyze performance at various levels of occlusion. While certain trackers can detect non-occluded objects well (over 35% AP), objects that are highly occluded, partially occluded, and out-of-frame remain challenging. Because of GTR’s overall strong performance in both detection and tracking, we use it as the basis for our subsequent experiments. We run all existing trackers at 1 fps and average AP across all categories with an IoU threshold of 0.5.
Detection Metrics Tracking Metrics
[…]
Baseline (GTR [64]) 0.78 13.24 37.54 14.18 28.19 16.02 8.86
Fine-tune entire model 0.52 10.36 24.08 10.34 17.93 7.70 3.93
Fine-tune regression head
+ proposal network
0.79 10.57 27.91 11.37 21.42 9.04 4.53
Fine-tune regression head 0.77 14.62 38.17 15.31 29.24 16.07 9.28
Amodal Expander 0.67 16.29 37.11 17.39 29.50 16.10 10.44 (+1.58)
Amodal Expander + ${}^{\dagger}$ 0.77 16.53 (+3.29) 37.80 (+0.26) 17.65 (+3.47) 29.96 (+1.77) 16.35 (+0.33) 10.28 (+1.42)
Table 3: Amodal expander on TAO-Amodal validation set. We ablate different strategies for repurposing GTR for amodal tracking, including fine-tuning all or part of the model, as well as our proposed Amodal Expander. The latter modestly outperforms naive fine-tuning of the regression head. When combined with additional augmentation techniques for generating synthetic occlusions such as PasteNOcclude (PnO) (as detailed in Sec 4.2), our final model produces noticeable gains for partially occluded and out-of-frame objects. All models (other than the baseline) were trained on TAO-Amodal training set for 20k iterations, while ${}^{\dagger}$ denotes 45k iterations of training. Table 4 focuses on performance for the people category, demonstrating even more impressive gains.
Detection Metrics Tracking Metrics
[…]
GTR [64] 0.29 37.15 71.49 42.07 53.81 17.47 14.39
FT regression head 0.41 49.32 78.93 53.26 61.36 20.44 18.74
Amodal Expander 2.26 71.64 84.07 73.74 74.22 26.77 (+9.30) 28.94
Amodal ${}^{\dagger}$ 2.46 (+2.17) 71.86 (+34.71) 84.21 (+12.72) 73.96 (+31.89) 74.34 (+20.53) 26.72 28.95 (+14.56)
Amodal Expander + PnO 1.94 69.87 83.86 72.58 73.20 26.68 28.76
Amodal Expander + ${}^{\dagger}$ 1.99 70.23 84.00 72.85 73.38 26.61 28.64
Table 4: Evaluating People on TAO-Amodal validation set. We follow the conventions of Table 3 but evaluate performance only on the people category. Here, our Amodal Expander improves over the baseline by more than 30% AP for partially occluded and out-of-frame people. This notable improvement could also be observed in the qualitative results in Figure 4. Training without PasteNOcclude (PnO) yields a slight performance drop. We posit that this dramatic performance increase comes from the fact that people is the most common category. Our results suggest that adding synthetic (occluded) examples is more helpful for less common categories.
4 Amodal tracking
Traditional and amodal tracking.
Given a sequence of images $I^{1},I^{2},...,I^{t}$, tracking approaches aims to output modal bounding boxes $b$, trajectory identifiers $\tau$, and class labels $s$ for objects across all frames. If an object is partially occluded, the box marks only the visible extent of the object, as illustrated in Fig. 1. We focus here on amodal trackers, which similarly takes as input a sequence of images, output modal boxes $b$, trajectory identifiers $\tau$, and classes $s$. Beyond this, amodal tracker further generates amodal boxes $b_{a}$, which cover full extent of partially and fully occluded objects.
In reality, training an entire amodal model is unfeasible due to the limited amount of amodal training data. Drawing inspiration from the human ability of amodal completion, we posit the feasibility of transforming a conventional tracker into an amodal one by leveraging its understanding of modal objects. We introduced a light-weight class-agnostic amodal expander $E$ in the next section.
4.1 Amodal expander
Amodal expander $E$ serves as a plug-in module to direct conventional trackers towards achieving amodal perception. For each object, amodal expander takes as input the modal box $b$ and features $f$, which can be retrieved from the tracker, and generates the amodal bounding boxes $b_{a}$.
Predicting amodal boxes in a residual manner.
In practice, the amodal expander is crafted in a residual manner, akin to the structure in two-stage detectors [46, 23] and trackers [64]. As illustrated in Fig. 2, the tracker initially produces region proposal features $f$ and modal proposals $b$, and subsequently refines $b$ through a regression head $R$ by predicting a modal box delta $\Delta b$. Our amodal expander receives modal box delta $\Delta b$ and object feature $f$ as input, generating amodal box delta. This delta is then applied to the modal proposal $b$ to generate amodal boxes $b_{a}$, denoted as $E(\Delta b,f)+b\approx b_{a}$. The training of the amodal expander basically follows the training of regression head [46] by matching box proposals with a ground truth and applying regression loss. We first match modal box predictions $b$ to a modal ground truth $b^{*}$. Then, we apply the regression loss, selected as smooth L1 [19], with the corresponding amodal ground truth $b^{*}_{a}$:
$L(b,\Delta b,f)=L_{reg}(E(\Delta b,f)+b,b^{*}_{a})$ (1)
As later shown in Table 5, the matching strategy is crucial in the training of the expander. Amodal expander architecture is introduced in Sec. 5.2. The effectiveness of amodal expander is demonstrated in Tabs. 3 and 4.
4.2 Crafting occlusion with PasteNOcclude (PnO)
To simulate occlusion scenarios during training, we introduced a data augmentation technique called PasteNOcclude (PnO). PasteNOcclude functions by integrating segments of objects into the original images to act as occludees. The segment collection comprises 505k objects extracted from LVIS [21] and COCO [38] images using segmentation masks. For each input image, we randomly select 1 to 7 segments from the collection and paste them at arbitrary locations, allowing for partial extension beyond the image boundary to replicate out-of-frame situations. Subsequently, we incorporate the ground truth boxes of the pasted segments into the original ground truth sets and apply regression loss according to the previously described pipeline. We find that PnO leads to improvements improvements in detection across all occlusion scenarios, shown in Table 3. We posit that this synthetic strategy is particularly important for the long-tailed nature of TAO-amodal, unlike COCO-amodal, where a similar synthetic occlusion strategy leads to limited improvement [65]. We provide visual examples of such synthetic occlusions in the appendix.
Figure 4: Qualitative results of Amodal Expander on TAO-Amodal val. Trackers augmented with our Amodal Expander produce predictions for both modal and amodal bounding boxes. In general, we are able to amodally complete objects that are occluded by objects in the scene (bottom-left) or objects that lie partially out of frame. Because we make use of online trackers that rely only on past evidence, we verify that our module can amodally complete objects that were occluded in the past as well as objects that become occluded later.
Detection AP Tracking AP
[…]
Modal 13.96 14.92 28.64 16.45 8.96
Amodal 16.41 17.64 29.87 16.35 10.13
Table 5: Ablation: Region proposal matching strategy. Given that existing off-the-shelf trackers generate modal proposals, direct matching with amodal ground truth yields suboptimal results. Our observation indicates that an improved strategy involves pairing region proposals with modal ground truth while applying regression loss to amodal predictions against the amodal ground truth. Both models are trained with PasteNOcclude (PnO) on TAO-Amodal training set for 20k iterations.
5 Empirical analysis
Our experiments are categorized into two parts. In Section 5.1, we assess the challenges of amodal detection and tracking by evaluating various off-the-shelf trackers. The efficacy of the amodal expander is examined in Section 5.2. Implementation details and ablation study could be found in the appendix.
5.1 Benchmarking state-of-the-art trackers
Evaluation Metrics.
Using the estimated visibility attributes, we assess the tracking and detection capabilities of the model through variations of detection AP [38] and Track-AP [10], representing the average precision across all categories at an IoU threshold of 0.5. We label objects with visibility less than 0.1 as heavily occluded, evaluated as AP[0.0, 0.1], where the superscript indicates the range of object visibility. If the visibility falls between 0.1 and 0.8, we categorize them as partially occluded, while those with visibility greater than 0.8 are considered non-occluded. Objects that stretch partially beyond the image boundary are referred to as out-of-frame (OoF) and evaluated with APOoF. Additionally, we assess the model’s performance on modal annotations with Modal AP. In tracking, we evaluate highly or partially occluded tracks (Track-AP[0, 0.8]), which are track with visibility at or below 0.8 for more than 5 frames (seconds). We also evaluate performance on modal annotations (Modal Track AP). A table is provided for easy reference to each metric definition in the appendix.
Evaluation on TAO-Amodal validation set.
We report detection and tracking average precision (AP) numbers of SOTA off-the-shelf trackers on TAO-Amodal validation set running at 1fps with an IoU threshold 0.5 in Tab. 2. We also observed similar performance trends when running at 5fps with higher IoU thresholds, shown in the appendix. Every off-the-shelf tracker was trained on either TAO [10] or LVIS [21], ensuring alignment of category vocabulary with our dataset. GTR [64] was trained on the combination of LVIS and COCO [38] by generating synthetic videos using the training strategy in [63]. ViTDet [37] was trained on LVIS and combined with online SORT [4] tracker. We reproduced AOA [12] using their released implementation, with object detector trained on LVIS and tracking ReID head trained on TAO. QDTrack [17] and TET [36] followed similar training procedures, pretraining detectors on LVIS and training instance similarity heads on TAO.
How well do standard trackers generalize to amodal perception?
Table 2 reveals notable differences in detection AP between modal (Modal AP) and amodal annotations (AP), amounting to an 8.49% difference. Additionally, the amodal tracking AP experiences a substantial decline compared to modal tracking AP. These findings confirm our expectation that existing trackers struggle to generalize effectively to amodal perception.
How well do standard trackers handle occlusion?
Existing off-the-shelf trackers exhibit reasonable performance in detecting non-occluded objects, with ViTDet achieving 39.71% AP[0.8,1] as revealed in Table 2. However, all trackers face challenges in detecting heavily occluded objects, and there is a noticeable decline in performance in partially occluded (AP[0.1,0.8]) and out-of-frame (OoF) scenarios. We noticed that ViTDet operating at 5 fps benefits from the property of SORT to estimate the location in the current frame using past information, resulting in improved results shown in the appendix. Nevertheless, this improvement comes at the cost of processing ViT-Det on 5x more frames than models running at 1 fps. In contrast, amodal completion could be a promising way for efficiently handling occlusion.
5.2 Amodal expander experiments
Amodal expander architecture could be referred in Fig. 3. We built the expander on top of GTR [64] as this method shows reasonable performance in both detection and tracking aspects in Table 2. The hidden dimension of MLP is selected as 256. ReLU [1] and dropout [48] with a probability of 0.2 were applied to each layer except the last one. The training of the amodal expander was conducted on the TAO-Amodal training set, along with PasteNOcclude (PnO) and augmentation used in [64]. All the modules except the amodal expander are frozen during training. Other ablation study, hyperparameter details for training and PnO could be found in the appendix.
Amodal expander or fine-tuning?
We evaluated amodal expander on TAO-Amodal validation set as shown in Tab. 3. Amodal expander trained with PnO for 45k iterations achieved 3.29% and 3.47% performance win under partially occluded (AP${}^{[0.1,0.8]}$) and out-of-frame (APOoF) scenario. We also experimented with different fine-tuning strategies. Fine-tuning entire model or solely the regression head and proposal network resulted in performance degradation. We posit that, with only 500 amodal training sequences, the models struggle to completely discard modal knowledge. Fine-tuning box regression head is suboptimal when compared to amodal expander. Amodal expander further provides flexibility to adjust the architecture and select different input information, which are both important as later shown in the ablation section.
Detecting people with amodal expander.
In Table 4, we study how well amodal expander detects and tracks people, which serves as a crucial category in many autonomous driving and tracking benchmarks. Amodal expander obtains a near 2x dramatic improvement compared to the baseline, with over 30% gain on AP[0.1, 0.8] and APOOF. Tracking on highly or partially occluded people (Track-AP[0.0,0.8]) are also increased by 14.56%. This shows that one can obtain a strong amodal people tracker that could also track objects of diverse category vocabulary with our dataset.
Proposal matching strategy matters.
To apply regression loss, training a box prediction head requires matching each region proposal to a ground truth box. A naive strategy is to directly match the region proposals to the amodal ground truth box. Nevertheless, as shown in Table 5, direct matching with amodal boxes leads to suboptimal results. We noted that since standard trackers generate modal region proposals, the model faced challenges in aligning proposals with the accurate ground truth due to a low Intersection over Union (IoU) between modal proposals and amodal ground truth. Consequently, we opted to match region proposals with modal bounding boxes and applied regression losses using amodal ground truth boxes.
6 Discussion
In this work, we focus on amodal perception of real-world objects. We draw inspiration from cognitive functions of amodal completion and object permanence in humans, that develop at an early age. Despite this, advancements in perception stacks like object detection and tracking, do not make amodal understanding central. We bring focus to three aspects/stages of building amodal perception stacks. First, we contribute a benchmark that annotates 880 categories of objects amodally in unconstrained indoor and outdoor settings, under partial and complete occlusion. Second, we contribute a benchmarking protocol in the form of metrics that evaluate detection and tracking specifically for the cases of partial or complete occlusions. Our key finding here is that existing algorithms fail dramatically on these metrics. To recover from this lack of performance on amodal detection and tracking and as a third contribution, we propose a light-weight and easy to train module, which we term as the Amodal Expander. Given modal boxes of objects, the expander learns to amodally complete the object bounding boxes till the extent of their occlusion. This simple fix gives us a boost of upto $\sim$30 AP points, specifically on the case of people.
Appendix
Detection AP0.5:${}_{0.5:0.95}$ Tracking AP0.5:${}_{0.5:0.95}$
@@ 6 Discussion
[…]
@@
Table 7: Off-the-shelf trackers on TAO-Amodal validation set running at 5 fps. In contrast to results shown in Table 2 in the main paper, ViTDet [37] with SORT [4] achieves a performance gain by running at a higher fps as SORT [4] leverages its capability to estimate the location in the current frame based on the location in previous frames. AP numbers are averaged across all categories at an IoU threshold 0.5.
@@
Table 8: Input to Amodal Expander. Modal box (deltas) $\Delta b$, output by the regression head as shown in Fig. 2 in the main paper, contains information about the exact location of modal box predictions. Object features $f$ encompass the visual appearance information of the modal proposals. We found that both information are important in amodally inferring the object’s shape. All models were trained on TAO-Amodal training set with PasteNOcclude (PnO) for 20k iterations.

### v3

@@ 3.1 Dataset statistics
We compare the statistics of TAO-Amodal to other amodal benchmarks in Table 1. For NuScenes, which only categorizes object visibilities into four buckets, we use interpolation to estimate the number of boxes below visibility 0.1 and 0.8. A few amodal datasets are omitted from the table, either because they have been incorporated into TAO-Amodal [7, 65] or because these datasets lack quantified visibilities for categorizing different occlusion scenarios [9, 54]. TAO-Amodal covers annotations across an extensive 833 categories, which can be used to learn and evaluate object priors in a large-vocabulary setting. Furthermore, TAO-Amodal features a 10$\times$ longer evaluation duration, ensuring a comprehensive evaluation.
3.2 Dataset Splits Design
Following TAO [10], we propose TAO-Amodal primarily as an evaluation benchmark. We choose to make the ‘validation’ and ‘test’ set larger to reliably benchmark trackers. We are not the first to do this: datasets in the single-object tracking [10] community have similarly focused on evaluation. With the success of foundation models trained on internet data, high quality evaluation benchmarks are more important than ever, as evidenced in the NLP community (e.g., MMLU [27]). Our training set is constructed in the spirit of instruction-tuning datasets: a small amount of data to align models to perform amodal tracking. We also find that increasing the size of the train set (by including test videos as train) does not significantly improve accuracy: adding 4×4$\times$4 × more training data only increases AP[0.1,0.8] from +3.3% to +3.7%, further validating our choice to dedicate limited annotated data for evaluation. See appendix for details.
Table 4: Multi-frame-aware amodal baselines on TAO-Amodal validation set. We explore extensions to include multi-frame signals for fine-tuned expander. Following [33], we use a Kalman filter to predict the positions of occluded objects, augmented by a monocular depth estimator to filter out spurious predictions. This leads to an increase in AP[0,0.1]. Further, we integrate multi-frame cross-attended Re-ID features, feeding them into the expander with concatenation. This boosts tracking and out-of-frame metrics.
Detection Metrics Tracking Metrics
[…]
Baseline (GTR [71]) 0.8 13.2 37.5 14.2 28.2 16.0 8.9
Amodal Expander 0.8 16.4 (+3.2) 37.7 (+0.2) 17.6 29.9 (+1.7) 16.4 10.1
+ Kalman filter 1.8 15.8 36.3 16.4 29.0 16.0 10.1
+ Depth [33] 2.0 (+1.2) 16.1 36.8 16.8 29.4 15.9 10.0
Amodal Expander
+ Temporal Re-ID
0.7 16.2 37.7 17.8 (+3.6) 29.8 17.1 (+1.1) 11.0 (+2.1)
4 Amodal tracking
4.0.1 Traditional and amodal tracking.
Given a sequence of images $I^{1},I^{2},...,I^{t}$, tracking approaches aim to output modal bounding boxes $b$, trajectory identifiers $\tau$, and class labels $s$ for objects across all frames. If an object is partially occluded, the box marks only the visible extent of the object, as illustrated in Fig. 2. We focus here on amodal trackers, which similarly take as input a sequence of images, but, in addition to the modal tracker outputs, they generate amodal boxes $b_{a}$, which cover the full extent of partially / fully occluded objects.
In practice, training an amodal tracker end-to-end is infeasible due to the limited amount of amodal training data. We focus instead on transforming a conventional tracker into an amodal one by leveraging its understanding of modal objects. To do this, we introduce a light-weight class-agnostic module $E$.
Table 5: Evaluating the ‘people’ category. We follow the conventions of Table 3 but evaluate performance only on the people category. Fine-tuned expander shows improvements over modal baseline, which can be observed in Fig. 7. We posit that this dramatic performance increase comes from the fact that people is the most common category. PasteNOcclude (PnO) leads to a slight drop for this category, which suggests that adding synthetic (occluded) examples is more helpful for less common categories.
Detection Metrics Tracking Metrics
[…]
GTR [71] 0.29 37.15 71.49 42.07 53.81 17.47 14.39
FT regression head 0.41 49.32 78.93 53.26 61.36 20.44 18.74
Amodal Expander 2.26 71.64 84.07 73.74 74.22 26.77 (+9.30) 28.94
Amodal Expander 2.46 (+2.17) 71.86 (+34.71) 84.21 (+12.72) 73.96 (+31.89) 74.34 (+20.53) 26.72 28.95 (+14.56)
Amodal Expander + PnO 1.94 69.87 83.86 72.58 73.20 26.68 28.76
Amodal Expander + PnO 1.99 70.23 84.00 72.85 73.38 26.61 28.64
4.1 Amodal expander
We design an amodal expander $E$, which serves as a plug-in module to conventional trackers. For each object, the amodal expander takes as input the modal box $b$ and an embedding $f$ (which can be extracted from the tracker), and generates amodal bounding boxes $b_{a}$.
4.1.1 Predicting amodal boxes in a residual manner.
The amodal expander operates as a refinement step, similar to the second stage of two-stage detectors [49, 25] and trackers [71]. We introduce the amodal expander in the context of GTR [71], although it can be applied to most standard modal trackers. As illustrated in Fig. 5, GTR produces modal boxes $b$ with corresponding object features $f$, and subsequently refines $b$ through a regression head $R$ by predicting a modal box delta $\Delta b$. Our amodal expander takes as input the modal box delta $\Delta b$ and object feature $f$ as input, generating an amodal box delta. This delta is then applied to the modal proposal $b$ to generate amodal boxes $b_{a}$, denoted as $E(\Delta b,f)+b\approx b_{a}$. The training of the amodal expander follows the training of regression head [49] by matching box proposals with a ground truth and applying regression loss. We first match modal box predictions $b$ to a modal ground truth $b^{*}$. Then, we apply the regression loss, selected as smooth L1 [21], with the corresponding amodal ground truth $b^{*}_{a}$:
$L(b,\Delta b,f)=L_{reg}(E(\Delta b,f)+b,b^{*}_{a})$ (1)
As shown in Tab. 6, the matching strategy is crucial for training the expander. Fig. 6 illustrates the amodal expander, and we provide implementation details in Sec. 5.2. We demonstrate the effectiveness of amodal expander in Tabs. 3 and 5.
Table 6: Ablation: Region proposal matching strategy. Given that modal trackers generate modal proposals, an improved strategy involves matching region proposals with modal ground truth (GT) while applying regression loss to amodal predictions against the amodal GT. Both expander models are trained with Paste-and-Occlude (PnO) on TAO-Amodal training set for 20k iterations.
Detection AP Tracking AP
[…]
Modal GT 13.96 14.92 28.64 16.45 8.96
Amodal GT 16.41 17.64 29.87 16.35 10.13
4.2 Synthesizing occlusion with Paste-and-Occlude (PnO)
To simulate occlusion scenarios during training, we use a data augmentation technique inspired by [72, 20], which we refer to as Paste-and-Occlude (PnO). Paste-and-Occlude functions by pasting object segments into the original images to act as occludees. The segment collection comprises 505k objects extracted from LVIS [23] and COCO [41] images using segmentation masks. For each input image, we randomly select 1 to 7 segments from the collection and paste them at arbitrary locations, allowing for partial extension beyond the image boundary to replicate out-of-frame situations. Subsequently, we incorporate the ground truth boxes of the pasted segments into the original ground truth masks. We find that PnO leads to improvements in detection across all occlusion scenarios, shown in Table 3. We posit that this synthetic strategy is particularly important for the long-tailed nature of TAO-amodal, unlike COCO-amodal, where a similar synthetic occlusion strategy leads to limited improvement [72]. We provide visual examples of synthetic occlusions in the appendix.
Figure 7: Qualitative results of Amodal Expander on TAO-Amodal val. Trackers fine-tuned with expander produce both modal and amodal predictions. The expander amodally complete objects that are occluded by objects in the scene (bottom-left) or objects that lie partially out of frame. We further verify that fine-tuned expander can amodally complete objects that were occluded in the past as well as objects that become occluded later.
5 Empirical analysis
In  Sec. 5.1, we assess the challenges of amodal detection and tracking by evaluating a number of amodal trackers and segmentors. Next, we investigate fine-tuning strategies and extensions for amodal baselines in Sec. 5.2. We present implementation details and further ablations in the appendix.
5.0.1 Evaluation Metrics.
Using the estimated visibility attributes, we assess the tracking and detection capabilities of the model through variations of detection AP [41] and Track-AP [10], representing the average precision across all categories at an IoU threshold of 0.5. We label objects with visibility less than 0.1 as heavily occluded, evaluated as AP[0.0, 0.1], where the superscript indicates the range of object visibility. If the visibility falls between 0.1 and 0.8, we categorize them as partially occluded, while those with visibility greater than 0.8 are considered non-occluded. Objects that extend beyond the image boundary are referred to as out-of-frame (OoF) and evaluated with APOoF. Additionally, we assess the model’s performance on modal annotations with Modal AP. In tracking, we evaluate highly or partially occluded tracks (Track-AP[0, 0.8]), which are track with visibility at or below 0.8 for more than 5 frames (seconds). We also evaluate performance on modal annotatiorns (Modal Track AP). We provide a table of metric definitions in the appendix for quick reference.
5.1 Benchmarking state-of-the-art trackers
Evaluation of modal detectors and trackers. We use three recent modal trackers, QDTrack [17], TET [39] and GTR [71] and a detector, ViTDet [40] for benchmarking. Every modal tracker is pre-trained on either TAO [10] or LVIS [23], ensuring alignment of category vocabulary with our dataset. GTR [71] is trained on the combination of LVIS and COCO [41] by generating synthetic videos using the training strategy in [70]. QDTrack [17] and TET [39] follow similar training procedures, pretraining detectors on LVIS and instance similarity heads on TAO for association. ViTDet [40] is trained on LVIS and combined with online SORT [4] tracker. We further adapt all models for amodal tracking by fine-tuning the regression head on TAO-Amodal training set for 20k iterations.
Evaluation of off-the-shelf amodal segmentors. While we finetune modal algorithms for the amodal task, we note that these may not be architecturally optimized for amodal perception. To this end, we also evaluate state-of-the-art methods from the amodal segmentation line of work (note that mask prediction is more prevalent for amodal perception than box prediction). We benchmark ORCNN [18], Amodal Mask-RCNN [18], AISFormer [58] and PCNet [67]. ORCNN proposes a loss which brings occluders and ocludees spatially close. Amodal Mask-RCNN trains an additional amodal mask head on top of Mask-RCNN. AISFormer, also based on Mask-RCNN, uses transformer blocks in its amodal mask head to learn the spatial relations between visible and occluded objects. These methods only need an image as input and are trained on COCOA-cls [18]. PCNet takes in modal masks of all objects in the scene as input, and recovers their relative ordering in the scene, before expanding modal masks into amodal ones. We use Detic [69] to get these modal masks. Finally, we run SORT [4] on top of all boxes obtained from aforementioned methods and evaluate with proposed metrics only on COCO classes. PCNet shines likely because it only needs to expand modal masks, which is a smaller lift than other baselines.
How well do SOTA methods handle amodal perception? In Tab. 2, we see that both amodal segmentation baselines and fine-tuned modal trackers struggle in handling heavy occlusion and out-of-frame cases. To bridge the gap, we further explore different fine-tuning schemes and effects of data augmentation in Tab. 3, introduced in the next section. We report the performance of modal trackers on TAO-Amodal validation set as an ablation in the appendix.
5.2 Building amodal baselines with amodal expander
We illustrate amodal expander architecture in Fig. 6. We build the expander on top of GTR [71] as this method shows reasonable performance in both detection and tracking aspects in Table 2, likely due to its transformer-based association architecture that links identities over longer time periods with a sliding window of size 16. The hidden dimension of MLP is 256. We apply ReLU [1] and dropout [51] with a probability of 0.2 to each layer except the last one. We train the amodal expander on the TAO-Amodal training set, along with PasteNOcclude (PnO) and augmentation used in GTR [71]. All the modules except the amodal expander are frozen during training. More ablation studies, hyperparameter details for training and PnO can be found in the appendix.
Explore fine-tuning strategies for amodal perception. We explored several fine-tuning strategies including amodal expander on TAO-Amodal validation set as shown in Tab. 3. Amodal expander trained with PnO for 45k iterations achieves 3.29% and 3.47% performance win under partially occluded (AP[0.1,0.8]) and out-of-frame (APOoF) scenario. Fine-tuning entire model or solely the regression head and proposal network results in performance degradation. We posit that, with only 500 amodal training sequences, the models struggle to completely discard modal knowledge. Fine-tuning box regression head is suboptimal when compared to amodal expander. Amodal expander further provides flexibility to adjust the architecture and select different input information, which are both important as shown in the ablation provided in appendix.
Integrating multi-frame signals into amodal baselines. In Tab. 4, we present two strategies for using temporal information within the amodal expander: 1) using a Kalman filter to forecast occluded object locations, with a monocular depth estimator to filter erroneous predictions, following [33], and 2) incorporating temporal Re-ID features. Note that (1) can associate single-frame detections, while also predicting new boxes when an object is completely occluded. This significantly improves AP[0,0.1]. For (2), we take multi-frame Re-ID features and feed them into the amodal expander with channel concatenation. This helps handle out-of-frame occlusion while improving tracking metrics.
Detecting people with amodal expander. In Table 5, we study how well the expander baseline detects and tracks people, which serves as a crucial category in many autonomous driving and tracking benchmarks. Amodal expander obtains a significant improvement compared to the modal baseline, particularly on AP[0.1, 0.8] and APOOF. Tracking on highly or partially occluded people (Track-AP[0.0,0.8]) also increases by 14.6%. This shows that one can obtain an effective amodal people tracker that could also track objects of diverse category vocabulary with our dataset using a simple fine-tuning scheme.
Importance of proposal matching strategies. To apply regression loss, training a box prediction head requires matching each region proposal to a ground truth box. A naive strategy is to directly match the region proposals to the amodal ground truth box. However, direct matching with amodal boxes leads to suboptimal results as shown in Tab. 6. As standard trackers generate modal region proposals, the model faced challenges in aligning proposals with the accurate ground truth due to a low Intersection over Union (IoU) between modal proposals and amodal ground truth. Matching proposals with modal boxes and applying regression loss using amodal ground truth yield better results.
6 Discussion
In this work, we focus on amodal perception of real-world objects. We draw inspiration from cognitive functions of amodal completion and object permanence in humans, that develop at an early age. Despite this, advancements in perception stacks like object detection and tracking, do not make amodal understanding central. We bring focus to three aspects/stages of building amodal perception stacks. First, we contribute a benchmark that annotates 833 categories of objects amodally in unconstrained indoor and outdoor settings, under partial and complete occlusion. Second, we contribute a benchmarking protocol in the form of metrics that evaluate detection and tracking specifically for the cases of partial or complete occlusions. Our key finding here is that existing algorithms struggle on these metrics. To address the observed limitations in amodal detection and tracking performance, and as a third contribution, we investigate data augmentation and fine-tuning schemes to boost existing tracking algorithms. We hope our empirical evaluation provides a foundation for improving amodal perception.
Appendix
Table 7: Off-the-shelf trackers on TAO-Amodal validation set. Off-the-shelf trackers were either trained on TAO [10] or on synthetic videos [71] generated using LVIS images [23], with categories aligned with our dataset. While certain trackers can detect non-occluded objects well (over 35% AP), objects that are highly occluded, partially occluded, and out-of-frame remain challenging, highlighting the difference between modal and amodal tracking. We run all existing trackers at 1 fps and average AP across all categories with an IoU threshold of 0.5.
@@ 6 Discussion
Table 8: Off-the-shelf trackers on TAO-Amodal validation with higher IoU thresholds. The definitions of our evaluation metrics can be found in Table 13. The AP numbers are averaged over 10 IoU values from 0.5 to 0.95 with a 0.05 step, denoted as AP0.5:0.95. We observed a similar performance trend as results evaluated with an IoU threshold 0.5. We run all trackers at 1 fps.
Detection AP0.5:0.95 Tracking AP0.5:0.95
@@
Table 9: Off-the-shelf trackers on TAO-Amodal validation set running at 5 fps. ViTDet [40] achieves a performance gain by running at a higher fps as SORT [4] leverages its capability to estimate the new location based on the location in previous frames. AP numbers are averaged across all categories at an IoU threshold 0.5.
@@
Table 10: Scaling up training data for amodal expander. All fine-tuning is done on a set of 1,928 videos, vs. 500 in the main paper.
Detection Metrics Tracking Metrics
[…]
Baseline (GTR [59]) 0.8 13.2 37.5 14.2 28.2 16.0 8.9
Fine-tune entire model 1.1 12.7 29.1 12.4 22.5 9.7 6.2
Fine-tune regression head 0.9 14.4 38.0 15.4 29.1 16.9 9.5
Amodal Expander 0.8 16.9 (+3.7) 37.7 17.9 (+3.7) 30.0 (+1.8) 16.5 10.7
Amodal Expander + PnO 0.7 16.5 37.8 17.9 (+3.7) 30.0 (+1.8) 16.5 10.8 (+1.9)
Table 11: Input to Amodal Expander. Modal box (deltas) $\Delta b$, output by the regression head as shown in  Fig. 5, contains information about the exact location of modal box predictions. Object features $f$ are embedded with visual appearance information of the modal proposals. We found that both information are important in amodally inferring the object’s shape. All models were trained on TAO-Amodal training set with PasteNOcclude (PnO) for 20k iterations.
@@
Table 12: Number of MLP layers in Amodal Expander. Empirically, a lightweight 2-layer MLP amodal expander is sufficient to generate reasonable amodal predictions. All models were trained on TAO-Amodal training set for 20k iterations.